I need help with this problem. Please explain step by how how to calculate on a calculator.

Assume a firm has earnings before depreciation and taxes of $550,000 and depreciation of $250,000.

a. If it is in a 30 percent tax bracket, compute its cash flow.

b. if it is in a 15 percent tax bracket, compute its cash flow.

c. cash flow benefits

The methods I used and answers I got were all wrong: (SEE BELOW)

If it is in a 30 percent tax bracket, compute its cash flow.

550,000 x (1-30%) = 385000

Cash Flow

385000

b. If it is in a 15 percent tax bracket, compute its cash flow:

(EBIT-Deprecation) x (1-30%) = (550,000-550,000) (1-0.15%) +550,000

Cash Flow

550,000

c. cash flow benefits = 550,000-385000 = 165000

Thanks!
Answer #1

a. If it is in a 30 percent tax bracket, compute its cash flow

=(earnings before depreciation and taxes-depreciation)*(1-tax rate)+depreciation

=(550000-250000)*(1-30%)+250000

=460000

b.

If it is in a 15 percent tax bracket, compute its cash flow

=(earnings before depreciation and taxes-depreciation)*(1-tax rate)+depreciation

=(550000-250000)*(1-15%)+250000

=505000

c. cash flow benefits

=505000-460000

=45000

the above is answer..